Consider the following information:

Imports $207.0
Net income from foreign investments 25.3
Foreign investments in U.S. 8.8
Government spending abroad 3.7
Exports 182.1
U.S. investments abroad 26.1
Foreign securities bought by U.S. 4.5
U.S. securities bought by foreigners 3.4
Purchase of short-term foreign securities 6.9
Foreign purchases of U.S. short-term securities 10.0

Determine the balance on the U.S. current account and capital accounts. Use a minus sign to enter the amount as a negative value. Round your answers to one decimal place.

Balance on current account: $  

Balance on capital account: $  
Answer #1

Current Account = Export - Import + Net Income from foreign investment =182.1-207.0+25.3=$0.4

Capital account= Investment by US in foreign - Investment in US by foreigners

=(3.7+26.1+4.5+6.9)-(8.8+3.4+10)=41.2-22.2=$19